The Keio Economic Observatory Review No. 8, March 1996

United States Labor Market Regulation^{∗}

Robert Evans Jr. Atran Professor of Labor Economics Brandeis University Visiting Professor Keio Economic Observatory Keio University and Abe Fellow

Traditionally among American economists the concept of regulation has been applied first to public utilities and secondarily to anti-trust issues under the Sherman and Clayton Anti-trust Acts Robinson-Patman Act, etc. The analysis has generally followed the perfectly competitive model which argues that the economy is at its most efficient when it operates with free markets. Regulations arise and views must be modified to recognize monopoly and other issues of market failure. Economists would grant that some regulations may be necessary, yet believe that regulation should take the form of market like instruments rather than bureaucratic rules when ever possible. An example of this in practice has been the introduction of saleable pollution rights as a substitute for the bureaucratic, or flat, rules which used to apply to all producers.

The wave of economic deregulation which has swept over parts of the United States has been centered in public utilities, electricity, telecommunications , etc. and in “public utility” type industries, airlines, trucking, etc. Before de-regulation many microeconomic studies of these industries had concluded that de-regulation would benefit the economy by giving rise to enhanced efficiency. This they argued would arise in one of two ways: a) inefficient operations would move to efficiency under conditions of increased competition,

b) the rents received by those who benefited from regulation would be reduced. Studies of specific industries following deregulation, for example, trucking and the airlines, have obtained empirical estimates of gains which were largely in accord with the predicted gains.

The American labor market is regulated in a number of ways but it is not a subject in which labor economists have taken a great deal of interest. The typical labor economics text book does not contain the word “regulation” in its index and neither does The Handbook of Labor Economics. Recently two books have been published which look at these issues: Christoph F. Bucechtemann, (Ed.), Employment Security and Labor Market Research, Ithaca: ILR Press, 1993 and Rebecca M. Blank, Social Protection and Economic Flexibility, Chicago: University of Chicago Press, 1994.

The results reported in those two volumes suggest that labor market regulations do not seem to have major distorting influences, though there are some questions as to whether the questions were adequately posed. For example, to test limitations on discharge and layoff they compared the speed of adjustments to economic shocks between countries with protections such as Germany and without, the United States. There were some differences, certainly in the degree to which adjustment came in hours or numbers of employed, but on the whole the authors concluded that the differences were not large. Yet the costs of regulation may not be measured in differential outcomes, but in other employer costs which were necessary in order to achieve the adjustments. In addition many of the authors subscribe to the idea that “workers are not bags of cement”, they therefore suggest that regulations which limit flexibility may encourage productivity and other “goods” in other ways, etc.

∗Reprinted and corrected from Keio Economic Observatory Occasional Paper E.No.16(March 1995).

Given the relationship between American unions and the Democratic Party in the United States and the long Democratic Party control of Congress, until this year, the policy thrust has been towards additional regulation. Consequently most analytical studies of labor market regulations in the United States don’t emphasize issues of deregulation. The focus of the analysis is on public policy aspects, whether the law has provided benefits to workers, how the regulations have been implemented, and an analysis of relevant court decisions. To a more limited extent there studies of the economic impact of the legislation or the regulation. Should there be movement toward deregulation in the American labor market, the methods of analysis which have estimated the impact of current regulations could be used to estimate the gain, and losses from deregulation.

If one envisions a simple supply and demand diagram for labor, many regulations can be visualized as forced shifts of the labor supply curve with in some cases changes in the elasticity over different portions of the supply curve. Clearly minimum wage laws and occupational licenses fall into this category. A few regulations, for example those in health and safety, may more properly be thought of as causing the demand curve to shift down. If, however, the safety standard is desired by employees their response would be to shift the supply curve down thereby paying part of the cost of the new standard. The various studies discussed below essentially try to estimate the shift in these curves. The efficiency loss to the economy could then be estimated in the traditional ways.

Briefly in paragraphs which follow the major regulations of the American labor market are discussed. These include which aspects have been studied, the methods of analysis and in a few cases some of the major conclusions. Certain generalities apply. Typically American economists do not generate their own data, rather they rely upon major government surveys for their data. Chief among these have been the National Longitudinal Surveys concerning the elderly and youth and the Current Population data which comes from the monthly labor force survey. The use of these large data sets have the obvious advantage that the sample size is quite large which allows for division along a number of attributes without loss of statistical significance. The disadvantage is that variables may not quite be what theory would dictate.

The method of analysis is to run regression equations of one form or another. Often the dependent variable will be “yes” or “no” making probit type regressions quite popular. The dependent variables of interest, wages, employment, the timing of events, etc. are regressed against an independent variable said to represent different regulatory states, and a series of control variables which hopefully hold other things constant. With the advent of changed levels of regulation it is sometime possible to engage in what the users like to refer to as a natural experiment. Thus before the new regulation takes place data are collected from specific firms or states and these data are compared with the same information after the advent of the new regulation. Alternatively before and after regression equations are run for supposedly similar states or firms some of which have been and the others have not been affected by the new regulations. The analytical tool is still the econometric equation but the independent variable is correctly specified and fewer control variables may have been necessary.

1. Labor Unions

The federal Wagner-Taft-Hartley laws and the decisions of the National Labor Relations Board as modified by the Federal Appeals Courts, and the Supreme Court provide protection for employees to organize into unions. In addition these laws and court decisions constrain employer behavior relative to unions and unionized workers. The coverage of these laws in terms of industries, size of firms, relationships to interstate commerce is quite broad. In addition to federal legislation a number of states have similar laws and regulatory bodies which extend coverage to additional employers. Executive orders at the federal level and state laws in many states have extended union protective legislation to government employees.

It is clear that legal support for unions has allowed their numbers to increase over what they would have been in the absence of legislation, and representation has given “voice” to workers in a way that would not have been possible without the legislation. The impact of legal protection for unions on the state of the economy is more difficult to estimate.

The decline in unionism between 1973 and 1986 has been explained by the union wage premium and the resulting growth in non-union employment within industries. The authors estimated that between 21 percent and 64 percent of the decline was explained depending upon the specification of the equations being estimated. The rather large difference in results depending upon the variables and their assumed relationship serves to underline how sensitive many conclusions are to the analytical methods employed.

A recent study looked at changes in employment, sales revenue, and investment in plant and equipment depending upon whether there was an election to choose a union in the company. The union election data came from the National labor Relations Board and the firm data from Standard and Poor’s Compustat [ This provides data on publicly trade firm based upon their annual reports and 10K reports. The latter are required by the Securities and Exchange Commission.] computer tape on firm characteristics. Except for a time period effect at the time of the election there was no effect on company behavior from having or not having a union. Other studies which have found equity losses following unionization, according to these authors, may reflect a transfer of income from owners to workers or (in my view) irrational stockholder behavior. Yet another study which used Bureau of Labor Statistics on union coverage and Compustat data found that industry specific effects explained about one-half of the variation in firm behavior, and that unions had strong negative effects on profitability working through investment and growth. Unions also limited employment growth and investments. Both of these studies with different results essentially ran regressions with the dependent variables being firm behavior, the independent variable union coverage with Compustat data to control for other differences in firm attributes.

The most recent union regulation has been the President’s just promised executive order prohibiting the federal government from buying goods from a firm which used permanent strike replacements. Whether the government can make such an end run around Supreme Court decisions is an interesting legal question. Its clear impact is to raise the cost of government services to the tax payer because government contracts are let on a low bid basis.

2. Anti-Discrimination

Beginning with the civil rights legislation in the 1960’s federal and state governments have extended rights of “equal treatment” to a variety of groups, based upon color, sex, national origin, religion, veterans’s status, , etc., and now most recently, physical and emotional handicaps. The first set of laws eliminated certain individual attributes from being the basis of employment, pay, and promotion decisions. In practice this has meant semi-quota regulations. The most recent require employers to not only ignore physical and emotional attributes, but also to make “reasonable” physical and other changes in order to accommodate the disabled worker. Most research is on whether firms comply or whether such rules have changed the ratio of the wages of disadvantaged individuals to advantaged individuals. This typically means tracking ratios over time before and after the regulations with suitable controls in an econometric equation or running separate regressions for say male and female wages and examining the coefficients to see how they differ. The cost to employers is seldom studied, partly because the rationale for the legislation is that such “disadvantaged” individuals are as productive as others. Thus there can only be minor costs, and to suggest otherwise leads to political problems. The emerging political battle over Affirmative Action in California is indicative of this.

3. Safety

Employers have a common law duty enforceable through the courts to compensate workers who are injured. Workers’ Compensation legislation was designed to provide compensation without long trials over fault, and with limitations on the level of individual recovery levels. Railway workers and seamen are treated outside these general laws. In addition at the federal level under the Occupational Health and Safety Administration and similar agencies in most states, employers are required to adhere to specific safety standards. Studies of OSHIA’s impact on safety have uniformly failed to find a relationship between its behavior and employee safety.

The main economic analysis concerning safety deals with cost-benefit studies. These studies ask whether the value of the lives or work days saved is greater than the cost to employers, and by implication to society of providing the protection. As is clear these studies require some agreement on the value of a saved life. Depending upon the approach used, compensating differentials, what the typical worker would pay to reduce the probability of accidental death at work, etc, a variety of different answers have been produced. Typically regulations are of the form, you must do “X” rather than injuries, lost time, or death rates must be improved to a given degree. For example in the environmental area, electric power companies were told they must have scrubbers in their smoke stacks, not that stack gas had to have less than “x” parts per liter. The former was quite expensive, but the latter, by switching the type of coal used was relatively cheap. Given the use of specific commands rather than the achievement of certain levels, there are a number of cost studies designed to see whether the same level of benefits could have been obtained more cheaply.

4. Wages under the Fair Labor Standards Act

Wage and Hour legislation provides for a federal minimum wage and for the payment of overtime at a rate of time and one-half after eight hours in a day or forty hours in a week. The latter applies to non-exempt employees which usually means non-supervisory. The U. S. Department of Labor estimates that the law applies to 30 percent of employers and 70 percent of employees. Some have estimated that non-compliance may reach 35 to 45 percent of all employment that should be subject to it. In addition to minimum wages there are federal and state prevailing wage laws which apply to federal and state government contracts. The most widely known one is the federal Davis Beacon Act. Recently the Commonwealth of Massachusetts attempted to eliminate via referendum its state “Davis Beacon” Act. The unions combined with other groups to defeat it.

The wage and hour laws also contain provisions concerning the employment of young people usually younger than 16 but sometimes extending to those who are in high school since there is clear evidence that working long hours and doing well in school don’t mix. There are, however, chicken and egg issues here because the poor progress in school my precede the long hours. In 1988 the Department of Labor estimated that 18 percent of young people were employed in violation of such regulations.

The minimum wage legislation is one of the more widely studied labor market regulations in the United States. One approach, termed by its some who use it, as a “natural experiment”, is to interview certain low wage employers such as fast food restaurants about their employment before the advent of a minimum wage change and then again afterward. Similarly some studies compare before and after the employment patterns for teenagers (assumed to be particularly affected by minimum wages) in various states using relative wages and proportions of impacted employment as state controls to study the impact of a uniform change in the federal minimum wage. Depending upon the researcher some recent studies have found very small employment changes or no employment impacts at all. Others have found that there are impacts. The results seem to depend upon the exact specification of the econometric equations which have been estimated. Various data sources have been used among them CPS data on individual and national longitudinal data for young workers.

Studies of the impacts upon employers in terms of profits or existence are seldom undertaken. This probably reflects both the nature of the political debate and the lack of an easy employer data source. The October 1992 issue of the Industrial and Labor Relations Review has good articles on the minimum wage. The April 1994 issue contains an exchange between authors who have looked at the same data and have come to different statistical results and conclusions.

5. Notice of Lay-Off

Under the Worker Adjustment and Retraining Notification Act of 1988 employers with 100 or more employees must provide 60 days notice of plant closure or a large layoff. When the law was being debated those in favor argued that with notice workers would be able to move more quickly to better jobs than if they did not have notice. Employer opposition was largely over the potential cost to them of not having enough workers at the end of the closing operations, and concerns over legislation moving into the delicate area of the regulation of employee discharge.

Evaluation of this law has essentially been limited to whether the “quick-er” and “better” has turned out to be true in practice. The analysis is the standard statistical one where the sample of those affected is drawn from a special labor force survey of displaced workers which is added to the Current Population Survey which consists of add on question to the monthly labor force survey on labor force participation and unemployment. Such studies allow one to distinguish between various levels of notice, and subsequent job finding, but usually lack an appropriate control group of individuals who have changed jobs without being characterized as displaced workers. Depending upon the type of worker and other factors there appear to be benefits to workers from the law. There is at least one study which has attempted to estimate the probability that an employee will leave if given notice and then relate this to the probability of being told. The data source is the same Displaced Worker Study, and statistical controls were used to provided the basis for an answer.

6. Employment at Will

Most workers in the United States are employed under the doctrine of employment at will, which means that an employer may discharge a worker for any reason at any time. Clearly for those individuals who belong to a union or have been protected under the anti-discrimination laws have protection against arbitrary discharge. In that sense employment at will has been severely compromised. Plaintiffs bar has also tried to carve out requirements of notice under employment manuals or alleged promises of employment which supposedly were relied upon by the now discharged workers. Legally there are three possible rationales for a court decision against the employer: a) a public policy exemption,

b) a decision that it is covered by a covenant of good faith, and c) That there is an issue of fair dealing. In 1980 13 states recognized one of these three and one state recognized all three. By 1989 45 states recognized at least 1 and 8 recognized all 3. As the number of individuals protected by one form of anti-discriminatory legislation grows, and more states recognize these exemptions, the potential for endless litigation increases. In many ways it is the uncertainly of result rather than the actual judgments that raise the economic cost of such laws to employers and ultimately to society. This is especially true in economic downturns, when the number of court cases grows significantly.

7. Immigration

The United States is entering a period when it has a greater number of immigrant households than it has had for many years. Foreign born individuals headed 7.9 percent of households in 1990 the highest level since 1940, though only about one-half the rates during the years 1890 to 1920.

Much of the economic analysis on immigration deals with the question of how well immigrants do in the U. S. labor market. The older studies suggested that correcting for individual characteristics immigrants originally earned about 17 percent less than natives but 30 years later they earned 11 percent more than natives. Recently these results have come under attack. A key fact has been that the more recent immigrants seem to have less human capital and human capital potential than did earlier generations of immigrants. Except for questions of welfare dependence, these results have little to do with whether levels or the nature of immigration should be controlled more than it now is.

In terms of control on rates of immigration the more relevant issue is the impact of immigration on the wages and earnings of the current labor force. One set of studies here has been to view large urban labor markets as closed systems, and then to estimate wage equations with a number of controls across these markets with the proportion of immigrants in each market as the dependent variable. These results suggest that a 10 percent increase in the proportion of immigrants would lead to 2 to 4 percent lower wages for the labor market. There also have been some studies of natural experiments, namely the large influx of Cubans a few years ago into Miami. There the influx did not seem to have any impact upon wage levels in that market.

City labor markets are not closed systems. To consider this issue requires the estimation of wage equations over time using the CPS data. This approach has suggested that one-third of the drop in wages for high school dropouts, 1980 to 1988, was due to the competition from new immigrants.

8. Social Security

There is an extensive literature on the fairness of Social Security between individuals with different work histories; on the impact of Social Security wealth and other plan characteristics on the decision to retire from the labor force; on the role of Social Security in the savings ratio, but no studies of which I am aware on the impact on the operation of the labor market, except labor force participation.

9. Private Pensions

No legalization requires employers to offer private pension plans which currently fall into three main types, Defined Benefit Plans, Defined Contribution Plans, and 401K Plans with defined benefit plans covering the largest number of employees. When employers offer pension plans they must conform to certain legal requirements. Two of the most important for defined benefit plans are vesting which must occur with in five years [Vesting means giving an employee the legal right to receive a pension even if the employment relationship is broken before the age of retirement] and payments to the Pension Guaranty Corporation which to a degree insures defined benefit plans.

10. Unemployment Insurance

This is a set of state program which were established in order to avoid a federal tax. No one expects these programs to be removed so the studies have been of the relationship between the amount of the wage replacement ratio and the length of time of unemployment, or the length of coverage and the timing of when employment is found. Finding employment tends to be bunched near the end of coverage suggesting that workers either wait, or they change their reservation wage. Similarly, increases in the replacement rate are said to cause workers to stay unemployed longer. In this latter case the analysis is based upon data in the National Longitudinal Survey and state administrative data. Since replacement rates will vary by states the analysis is a standard estimation of a labor force equation.

11. Employer Mandates

Almost any government requirement for private business may be considered a mandate. For example, some have asked whether the President’s new Minimum Wage proposal would run afoul of Congress’ new “no new unfunded mandates bill”. The answer is probably not since that law dealt with an impact on states. Minimum wages largely affect private firms though local governments are also covered and might now have to be made exempt. Yet, as a practical matter the use of the term “Employer Mandates” is usually reserved for very narrow government requirements. A classic example was 1978 federal legislation which required that employer health insurance packages had to contain comprehensive maternity insurance coverage. The nature of health insurance coverage has also been a popular state legislative requirement depending upon the power of your lobby [chiropractors] or your disease, [breast cancer].

There are only a few studies of mandates. Those associated with maternity benefits have been studies by “natural “experiments associated with the timing of pre-federal state mandates in some states and then the reverse when the federal law affected only those standards which had not gone first. The question there was whether wages of the benefiting group had gone down in order to pay for the mandated benefit. In this case the author argued that the answer was yes. The data source was that new American favorite the Current Population Survey and the analysis was to estimate wage equations with lots of controls.

12. Industrial Deregulation

The wage impact of industrial deregulation has been studied based upon the wage differentials by occupation between the regulated industries and other industries. It follows as a simple before and after analysis based upon special Bureau of Labor Statistics studies of some of these industries and the annual demographic file of the Current Population Survey.

13. Occupational Licenses

Approximately 500 different occupations from those such as medical doctors which require extensive education to others like hair dressing which do not, require licenses. Between 1950 and 1989 the number of workers in these licensed occupations grew from 3 percent to 18 percent of the labor force. Often in these occupations the quality of the output is difficult to determine or only becomes clear sometime after the service is purchased. It is the asymmetric aspect of the information which is said to lead to the need to license the producers. Yet if information is the problem, few if any occupations are set up to allow any one to practice so long as adequate information is provided. Within occupations one will occasionally observe such practices. A clear example is the Board Certification of Specialists in American medicine in which information is supplied but there are no limitations. Some hospitals, however, may limit privileges based upon board certification. There seems to be little interest in eliminating licenses, so there are few studies. One relatively recent one looked at rules governing optometry, essentially those which limited the growth of large chain optometrists. There the conclusion based upon comparisons of states with differing limitations was that the limits raised prices by 5 to 13 percent and did not affect quality. Other studies have found that average wages tend to be higher in those states with strict license requirements. Efforts to deregulate, if they can be called that, usually take the form of other occupations on the edge of the dominant one attempting to acquire certain of the dominant group’s rights. The rise of Chiropractors, the efforts of Nurse Practioners to obtain the right to practice independently and to prescribe medicine, and the recent effort through a successful referendum in the State of Washington of those who manufacture dentures to also fit them are examples.

14. Family Leave

The Family Leave Act requires employers to provide periods of unpaid leave so that workers can tend to medical problems of family members. Employers fear that it will be combined with other rules to make the impact more generous than originally conceived. It is all too new for their to have been any studies on the impact.

15. Employment Agencies

There is no law which requires employers to list jobs with the State Employment Agencies though there is a federal, and probably in a number of states, state executive orders which make this a requirement for those doing business with the federal or state government. I know of no studies directly examining their impact. I do not believe that studies of how individuals find jobs have shown any shift towards the Public Employment Office as a source. There also are private employment firms which provide the same services, for a fee, that the public Employment Agency does. These are often limited in the amount of money they may charge an individual and are subject to anti-discrimination laws.

Agencies which provide temporary employees [haken rodosha] are regulated, if at all, under state laws. To my knowledge there are few unique regulations for such agencies. There are now employee leasing firms which usually contract with an employer to supply the needed labor. In such cases there is continuity of workers at the job site but they are employees of the leasing firm. Certain economies of scale in personnel functions are the rational behind their use. Such arrangements raise issues of which firm is responsible for health and safety regulations, etc. States often pass laws to try and regulation these questions of agency. On the whole, none of these activities have been extensively studied.

A List of Sources

[1] Abraham, Katherine G, and Robert B. McKersie, (eds.), 1990, New Developments in the Labor Market, Cambridge: MIT Press.

[2] Addison, John T., and Pedro Portugal, 1992, “Advance Notice and Unemployment”, Industrial and Labor Relations Review 45 (July), pp. 645–664.

[3] —, Douglas A. Fox and Christopher J. Run, “The Impact of Advance Notice: A Comment on a Study by Nord and Ting”, Industrial and Labor Relations Review 45 (July), pp. 665–773.

[4] —, and McKinley L. Blackburn, 1994, “The Worker Adjustment and Retraining Notification Act”, Industrial and Labor Relations Review 47 (July), pp. 650–662.

[5] —, “Has WARN Warned?”, Journal of Labor Research 15 (Winter), pp. 83–90.

[6] Blank, Rebecca, M., (Ed.), 1994, Social Protection vs Economic Flexibility, Chicago: University of Chicago Press.

[7] Borjas, George J., 1994, “The Economics of Immigration”, Journal of Economic Literature 32 (Dec.), pp. 1667–1717.

[8] Bronars, Stephen G. and Donald R. Deere, 1990, “Union Representation Elections and Firm Profitability”, Industrial Relations 28 (Winter), pp. 15–37.

[9] —, 1993, “Union Organizing Activity, Firm Growth and the Business Cycle”, American Economic Review 83 (March), pp. 203–220.

[10] —, and Joseph Tracy, 1994, “The Effects of Unions on Firm Behavior”, Industrial Relations (October), 1994, pp. 426–451.

[11] Buechtemann, Christoph F. (Ed.), 1993, Employment Security and labor Market Behavior, Ithaca, New York: ILR Press.

[12] Card, David, 1990, “The Impact of the Mariel Boat lift on the Miami Labor Market”, Industrial and Labor Relations Review 43 (Jan.), pp. 245–257.

[13] —, 1992a, “Using Regional Variation in Wages to Measure the Effects of the Federal Minimum Wage”, Industrial and Labor Relations Review 46 (Oct), pp. 22–37.

[14] —, 1992b, “Do Minimum Wages Reduce Employment”, Industrial and Labor Relations Review 46 (Oct.), pp. 38–54.

[15] —, and Alan B. Krueger, 1994, “Minimum Wage and Employment”, American Economic Review 84 (Sept.), pp. 772–793.

[16] Cooter, Robert, 1994, “Market Affirmative Action”, San Diego Law Review (Winter), pp. 133–168.

[17] Edwards, Richard, 1993, Rights at Work, Washington: Brookings.

[18] Ehrenberg, Ronald G., 1994, Labor Markets and Integrating National Economies, Washington: Brookings.

[19] Employee Benefits Research Institute, 1987, Government mandates of Employee Benefits, Washington: EBRI.

[20] Fallick, Bruce Chelimsky, 1991, “Unemployment insurance and the Rate of Reemployment of Displaced Workers”, Review of Economics and Statistics 73 (May), pp. 228–235.

[21] Fishback, Price V. and Shaw Everett Kantor, 1994, “Did Workers Pay for the Passage of Workers’ Compensation Laws”, NBER Working Paper 4947.

[22] Gruber, Jonathan, 1994, “The incidence of Mandated Maternity Benefits”, American Economic Review 84 (June), pp. 622-641.

[23] Hass-Wilson, Deborah “The Effect of Commercial Practice Restrictions: The Case of Optometry”, Journal of Law and Economics 29 (April), pp. 165–186.

[24] Hendricks, Wallace, 1994, “Deregulation and Labor Earnings”, Journal of Labor Research 15 (Summer), pp. 207–234.

[25] Katz, Lawrence F. and Alan B. Krueger, 1992, “The Effects of the Minimum Wage on the Fast Food Industry”, Industrial and Labor Relations Review 46 (Oct), pp. 6–21.

[26] Krueger, Alan B., 1991, “The Evolution of Unjust Dismissal Legislation in the United States”, Industrial and Labor Relations Review (July), pp. 644–660.

[27] Krueger, Alan B., 1994, “The Effect of the Minimum Wage When It Really Bites”, NBER Working Paper 4757.

[28] McGarity, Thomas O., and Sidney A. Shapiro, 1993, Workers at Risk, New York: Praeger.

[29] Newmark, David, and William Watcher, 1992, “Employment Effects of Minimum and Sub minimum Wages”, Industrial and Labor Relations Review 46 (Oct.), pp. 55–81.

[30] Nord, Steven and Yuan Ting, 1991, “The Impact of Advance Notice of Plant Closing on Earnings and the Probability of Unemployment”, Industrial and Labor Relations Review 31 (July), pp. 681–691.

[31] Pearce, Thomas G., and James E. Groff, and John R. Wingender, 1995, “Union Decertification’s Impact on Shareholder Wealth”, Industrial Relations 34 (Jan.), pp. 58–72.

[32] Ruhm, Christopher J., 1994, “Advance Notice, Job Search, and Post Displacement Earnings”, Journal of Labor Economics 12 (Jan.), pp. 1–28.

[33] Ruser, John W. and Robert S. Smith, 1991, “Reestimating OSHA’s Effects”, Journal of Human Resources 26 (Spring), pp. 212–235.

[34] Wellington, Alison J., 1991, “Effects of the Minimum Wage on the Employment Status of Youths”, Journal of Human Resources 26 (Winter), pp. 27–46.

[35] Winston, Clifford, 1993, “Economic Deregulation”, Journal of Economic Literature 31 (Sept.), pp. 1263–1289.

[36] Viscus, W. Kip, 1992, Fatal Tradeoffs, Oxford: Oxford University Press.

[37] —,1993, “The Value of Risks to Life and Health”, Journal of Economic Literature 31 (Dec.), pp. 1212–1246.